# arXiv:2506.xxxxx

### v1

@@ Abstract
We propose a multi-field fiber inflation scenario in type IIB perturbative large volume compactifications, showing that the multi-field dynamics suppresses trans-Planckian displacements of the canonical inflaton. Considering a concrete K3-fibred Calabi-Yau (CY) threefold with $h^{1,1}({\rm CY})=3$ and having certain underlying symmetries, we show that the presence of multi-fibre moduli creates an assisted inflation scenario where multiple moduli collectively help in producing the cosmological observables consistent with the experimental observations. We further argue that individual field range excursions $(\Delta\phi_{n})$ corresponding to each of the inflaton fields can be estimated as $\Delta\phi_{n}=\Delta\phi/\sqrt{n}$, where $\Delta\phi$ denotes to the effective single-field inflaton range needed to generate the desired cosmological observables, and $n$ is the number of moduli assisting the multi-fibre inflation. We also present various numerical benchmark models consistently producing cosmological observables in the lights of the recent ACT experiments.
@@ 1 Introduction
Due to all these successful predictions, in the context of plain field theories, a plethora of inflationary models have been proposed over the past four decades. The simplest of those are single-field inflationary models which only depend on a few adjustable parameters. However, with increasingly precise bounds on so-called spectral index ($n_{s}$) and the tensor-to-scalar ratio ($r$) from improved cosmological observations, many minimal field-theoretic inflation models may be now excluded. Yet, even viable candidates in the context of plain field theory often fail to satisfy string-theoretic consistency requirements, particularly the swampland conjectures [4, 5] and the subsequent implications such as the refined de Sitter conjectures, issues related to trans-Planckian field ranges etc. (see for example [6, 7, 8, 9, 10] and reference therein). These constraints which are rooted in ultra-violet (UV) completion, severely limit the space of allowed inflationary potentials in string-derived EFTs. Furthermore, a crucial requirement for building viable low-energy effective string models -particularly those that incorporate inflation with a modulus serving as the inflaton- is the stabilization of all string moduli. Without stabilization, the moduli remain massless and dynamically unfixed, resulting in undetermined couplings, masses, and other physical parameters in the low-energy effective theory. This would render the theory phenomenologically and cosmologically inconsistent and unsuitable for making meaningful predictions. Consequently, a major task in string cosmology is to identify a moduli potential which, preferably exhibits a de Sitter vacuum and, upon minimization, consistently fixes the moduli at their vacuum expectation values. Furthermore, if the potential has the appropriate shape, one or more moduli can play the role of the inflaton field, thereby realizing the inflationary scenario. Some huge amount of efforts have been devoted to this pursuit over the past two decades, which have led to the development of various inflationary scenarios, e.g. see [11, 12] and references therein.
@@ 2.1.2 Perturbative LVS
With the inclusion of 1-loop effects– also known as the log-loop corrections – to the Kähler potential on top of the BBHL corrections used in the standard LVS, one arrives at an effectively modified overall volume ${\cal V}$ which we denote as ${\cal Y}$. It takes the following explicit form,
@@
where ${\cal Y}_{0}$ denotes the overall volume modified by $\alpha^{\prime}$ corrections appearing at string tree-level while ${\cal Y}_{1}$ is induced at string 1-loop level as given below [37, 79, 80, 38, 39, 81],
@@
Note that, the three parameters $\xi$, $\sigma$ and $\eta$ do not have any $g_{s}$ dependence which can be fixed by the $SL(2,\mathbb{Z})$ arguments [48], however $\sigma$ and $\eta$ can generically depend on the complex structure moduli. In order to keep track of this possibility we introduce $\sigma_{0}$ and $\eta_{0}$ in Eq. (2.13) as complex-structure moduli dependence parameters while still keeping the $SL(2,\mathbb{Z})$ motivated factors of $\zeta$ functions and the Euler characteristic of the CY. Subsequently we have the following pieces at leading order,
@@ 2.2 Sub-leading corrections to scalar potential
In the absence of any non-perturbative effects, there can be several types of perturbative effects which can induce useful scalar potential contributions, namely the BBHL’s $(\alpha^{\prime})^{3}$ corrections [19], the perturbative string-loop effects of [37] as well as the higher derivative F4 corrections of [30]. Using the Gukov-Vafa-Witten’s flux superpotential $W_{0}$ for stabilizing the complex-structure moduli and the axio-dilaton at their respective supersymmetric minimum, the perturbative scalar potential contributions can be collected in the following form
@@
Let us elaborate on each of these terms by giving some more details:
@@
  • The leading order contributions from the piece $V_{\alpha^{\prime}+{\rm log}\,g_{s}}$ appears at ${\cal O}({\cal V}^{-3})$ in the large volume expansion. This piece can be expressed as
@@
  • The so-called KK-type and Winding-type string loop effects are encoded as corrections to the Kähler potential, which in the Einstein frame reads as follows [31, 33, 34, 90, 35],
@@
  • Finally, the higher derivative F4-corrections cannot be captured by the conventional two-derivative corrections through the Kähler and superpotential. These corrections take the following simple form [30],
@@ 2.3 Ingredients for multi-field inflation
Here, the generic (non-flat) field space metric is represented as ${\cal G}_{ab}$ while ${\Gamma^{a}}_{bc}$’s denote the corresponding Christoffel connections, and the derivatives of the scalar potential are defined as: $\partial_{a}V=\frac{\partial V}{\partial\Phi^{a}}$. Further, the field space metric ${\cal G}_{ab}$ for the real moduli is obtained from the Kähler moduli space metric $K_{T_{\alpha}\overline{T}_{\beta}}$ using the following relation
@@
where $V_{\rm end}$ corresponds to the value of the potential at the end of inflation determined by $\epsilon=1$ and $m_{\rm inf}$ is the inflaton mass. From the single-field analysis, one usually has $N\simeq 50$ for Fibre inflation [57, 92, 93]. Following the constraints from the Planck 2018 data, one typically needs the scalar perturbation amplitude to satisfy $P_{s}\simeq 2.1\times 10^{-9}$ while the primordial scalar tilt is $n_{s}=0.9651\pm 0.0044$ and its running turns out to be $\alpha_{s}=-0.0041\pm 0.0067$ [71, 72]. In addition, the Atacama Cosmology Telescope (ACT) data gives $n_{s}=0.9666\pm 0.0077$, in agreement with Planck while a new constraint based on a combination of the Planck and ACT data (P-ACT) gives $n_{s}=0.9709\pm 0.0038$. Further, a combination of Planck, ACT, and DESI (P-ACT-LB) gives $n_{s}=0.9743\pm 0.0034$ and $\alpha_{s}=0.0062\pm 0.0052$ [73, 74, 75, 76].
@@ 3.1 An explicit CY threefold with $h^{1,1}({\rm CY})=3$
This volume form can also be expressed in the following form:
@@
where $Q_{\rm D_{3}}$ is twice the rhs of (3.31), for example see [23]. In fact, the relation (3.31) is motivated while looking at the connection the type IIB orientifold ingredients with those of the F-theory compactifications, which demands that rhs of 3.31 can take only integral values.
@@ 3.2 Various contributions to the scalar potential
In contrast to generic type IIB Calabi-Yau orientifold scenarios, our current construction exhibits an absence of several expected corrections. In particular,
@@
Further, notice that although this K3-fibred CY has several properties like a toroidal case, unlike the ${\mathbb{T}}^{4}$ case which has a vanishing (Chern numbers) $\Pi$, the K3 divisors have $\Pi=24$, and hence non-trivial higher derivative effects F4-corrections can be induced, taking the following form in the scalar potential,
@@
Summarizing all the contributions, we have [65]:
@@
In LVS setups, the overall volume ${\cal V}$ is fixed by the leading order contribution to the scalar potential and the full exchange symmetry seen from (3.38) is reduced to ${\mathbb{Z}}_{2}$, which helps in assisted fibre inflation with two fields as we will present in the upcoming section.
@@ 3.3 Single field fibre inflation
The minimal standard LVS scheme of moduli stabilisation fixes the CY volume ${\cal V}$ along with a small modulus $\tau_{s}$ controlling the volume of an exceptional del Pezzo divisor. Therefore any LVS models with 3 or more Kähler moduli, $h^{1,1}\geq 3$, can generically have flat directions at leading order. These flat directions are promising inflaton candidates with a potential generated at subleading order. Based on the geometric nature of the inflaton field and the source of inflaton potential, there are basically several classes of interesting inflationary models realized in the LVS framework. In this regard, fibre inflation happens to be one of the most popular inflationary models in LVS [54, 56, 57, 58]. The minimal design of fibre inflation consists of $3$-fields corresponding to the Kähler moduli of a K3-fibred CY threefold with one diagonal del-Pezzo divisor to support LVS. In this class of models, the typical volume form can be considered as ${\cal V}=\gamma_{b}\,\,\tau_{b}\,\sqrt{\tau_{f}}-\gamma_{s}\,\,\tau_{s}^{3/2}$ such that the overall volume ${\cal V}$ and small divisor volume $\tau_{s}$ are stabilized via the standard LVS prescription while the volume modulus corresponding to the K3-fibre $\tau_{f}$ drives the inflation through a sub-leading scalar potential effect encoded as 1-loop correction in the Kähler potential. It has been also observed that higher derivative F4-corrections can be useful for the purpose of realizing fibre-inflation like models, especially for the cases when KK-type string loop corrections are absent.
@@
where all the cosmological observables are evaluated at the horizon exit $\phi=\phi^{\ast}$. However, we also note that a combination of Planck, ACT, and DESI (P-ACT-LB) gives $n_{s}=0.9743\pm 0.0034$ and $\alpha_{s}=0.0062\pm 0.0052$ [73, 74, 75, 76]. Further, one typically has $N(\phi^{\ast})\gtrsim 60$ as an estimate for the sufficient $e$-folds needed to ensure a successful inflationary scenario. Further, the number of e-folding is typically determined by the following relation,
@@
This leads to the following tree level Kähler metric and it’s inverse metric,
@@
  • After fixing the overall volume ${\cal V}$, the quantities ${\cal R}_{1}\ll 1$ and ${\cal R}_{2}\ll 1$ serves as parameters which, as seen from (3.57), are further volume (${\cal V}$) suppressed as compared to ${\cal R}_{0}$. This is to be exploited in finding a sufficiently long plateau.
@@ 3.4 Two-field inflationary analysis
In this subsection we will present the two-field inflationary analysis for the benchmark models presented in the single-field case. For the single field analysis, we assumed that the overall volume modulus ${\cal V}$ sits at its perturbative LVS minimum which the $t^{2}$ modulus effectively drives the inflation. However, it is necessary to test the validity of such an assumption by considering the two-field analysis.
For this purpose, we consider the inflationary dynamics using the field basis $\Phi^{a}=\{{\cal V},t^{2}\}$ and will show that inflation is mostly driven by a single field while the overall volume modulus ${\cal V}$ remains seated in its perturbative LVS minimum. Now, the two-field scalar potential (3.36) takes the following form,
@@
where denotes derivatives w.r.t. the efolding $N$, i.e. ${\cal V}^{\prime}=\frac{d{\cal V}}{dN}$ etc. In addition, the $\epsilon$ parameter and the non-zero Christoffel connections, corresponding to the field space metric ${\cal G}_{ab}$ as defined in Eq. (3.55), are given as follows
@@ 4 Assisted Fibre Inflation in Perturbative LVS
From the detailed analysis presented in the previous section, we find that standard fibre inflation models typically need $\Delta\phi\simeq 6$Mp while the ACTivated ones require a relatively smaller values of the effective inflaton shifts for successfully realizing the cosmological observables. In this section we present the three-field evolution analysis, where the overall volume ${\cal V}$ remains at the perturbative LVS minimum while the other two modulus assist in driving the multi-field fibre inflation.
@@ 4.1 Ingredients for three-field evolution
In our approach, we adopt the real field basis $\Phi^{a}={{\cal V},t^{2},t^{3}}$ and utilize the CY threefold?s symmetry so that $t^{2}$ and $t^{3}$ drive inflation, while ${\cal V}$ stays largely fixed at its perturbative LVS minimum. Using the leading order tree-level Kähler metric (3.25), the field space metric and its inverse metric are given by,
@@
Here $\phi^{a\ast}$ corresponds to the horizon exit while $\langle\phi^{a}\rangle$ denotes the moduli VEVs at the LVS minimum. As argued earlier, for any choice of canonical field basis, the inflaton shift $\Delta\phi$ remains invariant. However, the individual fields may result in different shifts in a different choice of basis. For example, let us demonstrate the argument by considering three choice of canonical bases relevant for different purposes.
@@
    This canonical basis $\varphi^{a}$ respects the exchange symmetry $1\leftrightarrow 2\leftrightarrow 3$, however it does not have the overall volume ${\cal V}$ as a singled our direction. This is needed for manifestation of the mass hierarchy between the overall volume and the remaining two volume moduli.
@@
    $\hskip-28.45274pt{\cal V}=\frac{1}{\sqrt{2}}e^{\sqrt{\frac{3}{2}}\phi^{1}},\quad t^{2}=\frac{1}{\sqrt{2}}e^{\frac{1}{\sqrt{6}}\left(\phi^{1}-\left(1+{\sqrt{3}}\right)\phi^{2}+\left(1+{\sqrt{3}}\right)\right)\phi^{3}},$ (4.21)
    $\hskip-28.45274ptt^{3}=\frac{1}{\sqrt{2}}e^{\frac{1}{\sqrt{6}}\left(\phi^{1}+\left(1+{\sqrt{3}}\right)\phi^{2}+\left(1-{\sqrt{3}}\right)\right)\phi^{3}}.$

### v2

# arXiv: 2506.22630

@@ Abstract
We propose a multi-field fibre inflation scenario in type IIB perturbative large volume compactifications, showing that the multi-field dynamics suppresses trans-Planckian displacements of the canonical inflaton. Considering a concrete K3-fibred Calabi-Yau (CY) threefold with $h^{1,1}({\rm CY})=3$ and having certain underlying symmetries, we show that the presence of multi-fibre moduli creates an assisted inflation scenario where multiple moduli collectively help in producing the cosmological observables consistent with the current experimental bounds. We further argue that individual field range excursions $(\Delta\phi_{n})$ corresponding to each of the inflaton fields can be estimated as $\Delta\phi_{n}=\Delta\phi/\sqrt{n}$, where $\Delta\phi$ denotes the effective single-field inflaton range needed to generate the desired cosmological observables, and $n$ is the number of moduli assisting the multi-fibre inflation. We also present various numerical benchmark models consistently producing cosmological observables in light of the recent ACT experiments.
@@ 1 Introduction
Due to all these successful predictions, in the context of plain field theories, a plethora of inflationary models have been proposed over the past four decades. The simplest of those are single-field inflationary models which only depend on a few adjustable parameters. However, with increasingly precise bounds on so-called spectral index ($n_{s}$) and the tensor-to-scalar ratio ($r$) from improved cosmological observations, many minimal field-theoretic inflation models may be now excluded. Yet, even viable candidates in the context of plain field theory often fail to satisfy string-theoretic consistency requirements, particularly the swampland conjectures [4, 5] and the subsequent implications such as the (refined) de Sitter conjectures [6, 7], issues related to trans-Planckian field ranges [8, 9, 10, 11, 12] etc. (for review see [13, 14] and reference therein). These constraints which are rooted in ultra-violet (UV) completion, severely limit the space of allowed inflationary potentials in string-derived EFTs. Furthermore, a crucial requirement for building viable low-energy effective string models -particularly those that incorporate inflation with a modulus serving as the inflaton- is the stabilization of all string moduli. Without stabilization, the moduli remain massless and dynamically unfixed, resulting in undetermined couplings, masses, and other physical parameters in the low-energy effective theory. This would render the theory phenomenologically and cosmologically inconsistent and unsuitable for making meaningful predictions. Consequently, a major task in string cosmology is to identify a moduli potential which, preferably exhibits a de Sitter vacuum and, upon minimization, consistently fixes the moduli at their vacuum expectation values. Furthermore, if the potential has the appropriate shape, one or more moduli can play the role of the inflaton field, thereby realizing the inflationary scenario. Some huge amount of efforts have been devoted to this pursuit over the past two decades, which have led to the development of various inflationary scenarios, e.g. see [15, 16] and references therein.
@@ 2.1.2 Perturbative LVS
With the inclusion of one-loop effects –also known as log-loop corrections– to the Kähler potential on top of the BBHL corrections used in the standard LVS, one arrives at an effectively modified overall volume ${\cal V}$ which we denote as ${\cal Y}$. It takes the following explicit form,
@@
where ${\cal Y}_{0}$ denotes the overall volume modified by $\alpha^{\prime}$ corrections appearing at string tree-level while ${\cal Y}_{1}$ is induced at string one-loop level as given below [41, 83, 84, 42, 43, 85],
@@
Note that, the three parameters $\xi$, $\sigma$ and $\eta$ do not have any $g_{s}$ dependence which can be fixed by the $SL(2,\mathbb{Z})$ arguments [52], however $\sigma$ and $\eta$ can generically depend on the complex structure moduli. In order to keep track of this possibility we introduce $\sigma_{0}$ and $\eta_{0}$ in Eq. (2.13) as complex-structure moduli dependence parameters while still keeping the $SL(2,\mathbb{Z})$ motivated factors of the Riemann $\zeta$-functions and the Euler characteristic of the CY. Subsequently we have the following pieces at leading order,
@@ 2.2 Sub-leading corrections to scalar potential
In the absence of any non-perturbative effects, there can be several types of perturbative effects which can induce useful scalar potential contributions, namely the BBHL’s $(\alpha^{\prime})^{3}$ corrections [23], the perturbative string-loop effects of [41] as well as the higher derivative F4 corrections of [34]. Using the Gukov-Vafa-Witten’s (GVW) flux superpotential $W_{0}$ for stabilizing the complex-structure moduli and the axio-dilaton at their respective supersymmetric minimum, the perturbative scalar potential contributions can be collected in the following form
@@
We now examine each term in greater detail.
@@
  • The leading order contributions from the piece $V_{\alpha^{\prime}+{\rm log}\,g_{s}}$ appears at ${\cal O}({\cal V}^{-3})$ in the large volume expansion, which can be expressed as follows
@@
  • The so-called KK-type and Winding-type string loop effects are encoded as corrections to the Kähler potential, which in the Einstein frame they are written as follows [35, 37, 38, 94, 39],
@@
  • Finally, the higher derivative F4-corrections cannot be captured by the conventional two-derivative corrections through the Kähler potential and the superpotential. These corrections take the following simple form [34],
@@ 2.3 Ingredients for multi-field inflation
Here, the generic (non-flat) field space metric is represented as ${\cal G}_{ab}$ while ${\Gamma^{a}}_{bc}$’s denote the corresponding Christoffel connections, and the derivatives of the scalar potential are defined as: $\partial_{a}V=\frac{\partial V}{\partial\Phi^{a}}$. Furthermore, the field space metric ${\cal G}_{ab}$ for the real moduli is obtained from the Kähler moduli space metric $K_{T_{\alpha}\overline{T}_{\beta}}$ using the following relation
@@
where $V_{\rm end}$ corresponds to the value of the potential at the end of inflation determined by $\epsilon=1$ and $m_{\rm inf}$ is the inflaton mass. From the single-field analysis, one usually has $N\simeq 50$ for Fibre inflation [61, 96, 97]. Following the constraints from the Planck 2018 data, one typically needs the scalar perturbation amplitude to satisfy $P_{s}\simeq 2.1\times 10^{-9}$ while the primordial scalar tilt is $n_{s}=0.9651\pm 0.0044$ and its running turns out to be $\alpha_{s}=-0.0041\pm 0.0067$ [75, 76]. In addition, the Atacama Cosmology Telescope (ACT) data gives $n_{s}=0.9666\pm 0.0077$, in agreement with Planck while a new constraint based on a combination of the Planck and ACT data (P-ACT) gives $n_{s}=0.9709\pm 0.0038$. Furthermore, a combination of Planck, ACT, and DESI (P-ACT-LB) gives $n_{s}=0.9743\pm 0.0034$ and $\alpha_{s}=0.0062\pm 0.0052$ [77, 78, 79, 80].
@@ 3.1 An explicit CY threefold with $h^{1,1}({\rm CY})=3$
Using the above relations between the four- and two-cycle moduli, the volume can also be expressed in the following form:
@@
where $Q_{\rm D_{3}}$ is twice the right-hand-side (rhs) of (3.31), for example see [27]. In fact, the relation (3.31) is motivated while looking at the connection the type IIB orientifold ingredients with those of the F-theory compactifications, which demands that rhs of 3.31 can take only integral values.
@@ 3.2 Various contributions to the scalar potential
Unlike generic type IIB Calabi-Yau orientifold scenarios, our current construction features a vanishing of several expected corrections. In particular,
@@
Moreover, notice that although this K3-fibred CY has several properties like a toroidal case, unlike the ${\mathbb{T}}^{4}$ case which has a vanishing $\Pi$ (Chern numbers), the K3 divisors have $\Pi=24$, and hence non-trivial higher derivative effects F4-corrections can be induced, taking the following form in the scalar potential,
@@
Collecting all the contributions, we have the following scalar potential [69]:
@@
In LVS frameworks, the overall volume ${\cal V}$ is fixed by the leading order contribution to the scalar potential and the full exchange symmetry seen from (3.38) is reduced to ${\mathbb{Z}}_{2}$, which helps in assisted fibre inflation with two fields as we will present in the upcoming section.
@@ 3.3 Single field fibre inflation
The minimal standard LVS scheme of moduli stabilisation stabilizes the CY volume ${\cal V}$ along with a small modulus $\tau_{s}$ controlling the volume of an exceptional del Pezzo divisor. Therefore any LVS models with 3 or more Kähler moduli, $h^{1,1}\geq 3$, can generically have flat directions at leading order. These flat directions are promising inflaton candidates with a potential generated at subleading order. Based on the geometric nature of the inflaton field and the source of inflaton potential, there are basically several classes of interesting inflationary models realized in the LVS framework. In this regard, fibre inflation happens to be one of the most popular inflationary models in LVS [58, 60, 61, 62]. The minimal design of fibre inflation consists of $3$-fields corresponding to the Kähler moduli of a K3-fibred CY threefold with one diagonal del-Pezzo divisor to support LVS. In this class of models, the typical volume form can be considered as ${\cal V}=\gamma_{b}\,\,\tau_{b}\,\sqrt{\tau_{f}}-\gamma_{s}\,\,\tau_{s}^{3/2}$ such that the overall volume ${\cal V}$ and small divisor volume $\tau_{s}$ are stabilized via the standard LVS prescription while the volume modulus corresponding to the K3-fibre $\tau_{f}$ drives the inflation through a sub-leading scalar potential effect encoded as one-loop correction in the Kähler potential. It has been also observed that higher derivative F4-corrections can be useful for the purpose of realizing fibre-inflation like models, especially for the cases when KK-type string loop corrections are absent.
@@
where all the cosmological observables are evaluated at the horizon exit $\phi=\phi^{\ast}$. However, we also note that a combination of Planck, ACT, and DESI (P-ACT-LB) gives $n_{s}=0.9743\pm 0.0034$ and $\alpha_{s}=0.0062\pm 0.0052$ [77, 78, 79, 80]. Further, one typically has $N(\phi^{\ast})\gtrsim 60$ as an estimate for the sufficient $e$-folds needed to ensure a successful inflationary scenario. Moreover, the number of e-folding is typically determined by the following relation,
@@
This leads to the following tree-level Kähler metric and its inverse metric,
@@
  • After fixing the overall volume ${\cal V}$, the quantities ${\cal R}_{1}\ll 1$ and ${\cal R}_{2}\ll 1$ serve as parameters which, as seen from (3.57), are further volume (${\cal V}$) suppressed as compared to ${\cal R}_{0}$. This is to be exploited in finding a sufficiently long plateau.
@@ 3.4 Two-field inflationary analysis
In this subsection, we will present the two-field inflationary analysis for the benchmark models presented in the single-field case. For the single-field analysis, we assumed that the overall volume modulus ${\cal V}$ sits at its perturbative LVS minimum while the $t^{2}$ modulus effectively drives the inflation. However, it is necessary to test the validity of such an assumption by considering a two-field analysis.
For this purpose, we consider the inflationary dynamics using the field basis $\Phi^{a}=\{{\cal V},t^{2}\}$ and show that inflation is mostly driven by a single field while the overall volume modulus ${\cal V}$ remains seated in its perturbative LVS minimum. Now imposing $t^{2}=t^{3}$ in the scalar potential (3.36), the effective two-field potential takes the following form,
@@
where denotes derivatives w.r.t. the $e$-folding $N$, i.e. ${\cal V}^{\prime}=\frac{d{\cal V}}{dN}$ etc. In addition, the $\epsilon$ parameter and the non-zero Christoffel connections, corresponding to the field space metric ${\cal G}_{ab}$ as defined in Eq. (3.55), are given as follows
@@ 4 Assisted Fibre Inflation in Perturbative LVS
From the detailed analysis presented in the previous section, we find that standard fibre inflation models typically need $\Delta\phi\simeq 6$Mp while the ACTivated ones require relatively smaller values of the effective inflaton shifts for successfully realizing the cosmological observables. In this section we present the three-field evolution analysis, where the overall volume ${\cal V}$ remains at the perturbative LVS minimum while the other two moduli assist in driving the multi-field fibre inflation.
@@ 4.1 Ingredients for three-field evolution
In our approach, we adopt the real field basis $\Phi^{a}=\{{\cal V},t^{2},t^{3}\}$ and employ the symmetry of the CY threefold so that $t^{2}$ and $t^{3}$ drive inflation, while ${\cal V}$ stays largely fixed at its perturbative LVS minimum. Using the leading-order tree-level Kähler metric (3.25), the field space metric and its inverse metric are given by,
@@
Here $\phi^{a\ast}$ corresponds to the horizon exit while $\langle\phi^{a}\rangle$ denotes the moduli VEVs at the LVS minimum. As argued earlier, for any choice of canonical field basis, the inflaton shift $\Delta\phi$ remains invariant. However, the individual fields may result in different shifts in a different choice of basis. To exemplify this, we analyze three canonical bases, each serving a distinct purpose.
@@
    This canonical basis $\varphi^{a}$ respects the exchange symmetry $1\leftrightarrow 2\leftrightarrow 3$, but it does not treat the overall volume $\mathcal{V}$ as a singled-out direction. This is required to establish the mass hierarchy between the overall volume modulus and the other two volume moduli.
@@
    $\hskip-28.45274pt{\cal V}=\frac{1}{\sqrt{2}}e^{\sqrt{\frac{3}{2}}\phi^{1}},\quad t^{2}=\frac{1}{\sqrt{2}}e^{\frac{1}{\sqrt{6}}\left(\phi^{1}-\left(1-{\sqrt{3}}\right)\phi^{2}+\left(1+{\sqrt{3}}\right)\right)\phi^{3}},$ (4.21)
    $\hskip-28.45274ptt^{3}=\frac{1}{\sqrt{2}}e^{\frac{1}{\sqrt{6}}\left(\phi^{1}+\left(1+{\sqrt{3}}\right)\phi^{2}-\left(1-{\sqrt{3}}\right)\right)\phi^{3}}.$